OpenStudy (anonymous):

Consider the sequence 16, –8, 4, –2, 1, ... a. Describe the pattern formed in the sequence. b. Find the next three terms.

ganeshie8 (ganeshie8):

do u see any pattern ?

ganeshie8 (ganeshie8):

16, -8, 4, -2, 1.... 16 --- = ? -2

OpenStudy (anonymous):

yeah

ganeshie8 (ganeshie8):

wat is it ? :)

OpenStudy (anonymous):

8

ganeshie8 (ganeshie8):

we can get next number by multiplying -1/2, right ?

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

16 * (-1/2) = -8 -8 * (-1/2) = 4 4*(-1/2) = -2 -2*(-1/2) = 1 1*(-1/2) = ?